OpenStudy (anonymous):

What is the simplified form of 3 over x cubed all over 1 over x squared ?

OpenStudy (anonymous):

3/x

OpenStudy (kropot72):

\[\large \frac{\frac{3}{x ^{3}}}{\frac{1}{x ^{2}}}=\frac{3}{x ^{3}} \times \frac{x ^{2}}{1}\]